# Full-sky Models of Galactic Microwave Emission and Polarization at Sub-arcminute Scales for the Python Sky Model

### v1

@@ 4.2 CO Models
co1 Single line emissions at 115, 230, 346 GHz Planck PR2 Type-1 maps smoothed to $1^{\circ}$
co2 "+ $0.1\%$ polarized "
co3 " " Simulated high galactic clouds
@@
Galactic CO line emission is strong enough that the $J=1\rightarrow 0$, $J=2\rightarrow 1$, $J=3\rightarrow 2$ transitions have been detected even in the broad photometric Planck bands (Planck Collaboration XIII, 2014; Planck Collaboration X, 2016). CO emission can be polarized (Goldreich & Kylafis, 1981) and so may be a relevant foreground for CMB polarization analyses (Puglisi et al., 2017). In this section, we describe the implementation of three CO models into the PySM 3 framework. These models are summarized in Table 5.
Because of the intrinsically low degree of CO polarization and the long integration time required to achieve a significant detection, it is difficult to carry out CO polarization surveys, and wide area measurements have not yet been made. In the absence of data-based templates, the approach to modeling CO has been to assume a small degree of polarization applied to total intensity maps. For instance, Puglisi et al. (2017) presented a model to simulate the polarized emission of CO lines in molecular clouds at high Galactic latitudes by considering the 3D spatial distribution of CO in the Galaxy.
All CO models implemented here are based on the CO $J=1\rightarrow 0$, $J=2\rightarrow 1$, $J=3\rightarrow 2$ Type-1 intensity maps [^12] HFI_CompMap_CO-Type1_2048_R2.00.fits derived by Planck Collaboration XIII (2014) from Planck data. The CO maps are obtained exploiting the mismatches in the detector bandpasses to recover the CO from the CMB and the other Galactic foregrounds using the Modified Internal Linear Combination Algorithm (MILCA). The templates produced from this analysis are preferred to direct measurements from existing spectroscopic surveys (e.g., Dame et al., 2001) because they are not limited to low Galactic latitudes and have a uniform angular resolution over the sky.
Planck Collaboration XIII (2014) also produced Type-2 templates based on a multi-frequency analysis. However, this method is more prone to modeling errors in separating the Galactic dust emission from the CO emission, and so we prefer the Type-1 templates for our purposes.
@@
The second CO model co2 introduces a simple implementation of polarization. The $I$ maps of the three transitions used in co1 are converted to $P$ maps using a sky-constant, user-defined polarization fraction $p$, where the default is $p=0.1$%. The $Q$ and $U$ maps are made from the $P$ maps using the polarization angle of the thermal dust emission as determined by a component separation analysis with Commander (Planck Collaboration X, 2016).
Finally, the third CO model co3 not only accounts for the polarized emission as in co2 but also includes small-scale emission from high-Galactic latitude clouds. We perform a dedicated simulation with the LogSpiral model from Puglisi et al. (2017), which provided the best fit to Planck data, to produce maps of the contribution to the total and polarized CO intensity at sub-degree scales from molecular clouds at high Galactic latitudes. The $I$, $Q$, and $U$ maps generated by the simulation are then added to the $I$, $Q$, and $U$ maps of co2 to produce the final co3 model. This injection introduces CO emission at scales $\lesssim 1^{\circ}$ in regions where the CO templates have very low signal-to-noise ratio.
5 Validation and Characterization of Models
In this section, we compare the new foreground models presented in this work both to data and to previous models. In Section 5.1, map-based comparisons are made to visually highlight differences between the models and observational data. In Section 5.2, we compare power spectra of the dust and synchrotron models with observations. After developing the methodology in Section 5.2.1, we demonstrate that the two-point statistics of the stochastic small scales are properly modulated for different regions of sky defined by Galactic masks of varying size. Section 5.2.2 compares the dust models to Planck NPIPE maps at 353 GHz. Section 5.2.3 compares the PySM synchrotron models with the synchrotron map from the BeyondPlanck analysis for intensity, and with the Planck Revisited reprocessed maps for polarization. Additional validation is performed in the sky patch observed by the BICEP/Keck telescopes in Section 5.2.4. Section 5.3 compares the level of frequency decorrelation in dust $BB$ emission across the models, demonstrating that all models respect current constraints from Planck observations. Section 5.4 quantifies the level of extragalactic contamination in the maps. In Section 5.5, we assess the level of non-Gaussianity introduced by the polarization fraction tensor formalism.
@@ 5 Validation and Characterization of Models
5.1 Maps
We provide several map-level comparisons between PySM dust models and observations. Specifically, we compare data from the Planck third data release PR3 Planck Collaboration III (2020) and PySM dust models d1, which was widely used as a reference in previous versions of the PySM, and is described in Thorne et al. (2017), d9, which is similar to d1 but uses different input templates and includes small-scale fluctuations of template maps and spectral parameters, as described in Sec. 3, and d12, which is conceptually quite different from the other two models. The goal is to provide a qualitative comparison of the spatial characteristics of the models at small scales. We select $16.7^{\circ}\times 16.7^{\circ}$ patches of the sky, and compare the data and models at 353 GHz in both total and polarized intensity. We focus on two patches, one close to the Galactic plane with $(l,b)=(180^{\circ},-10^{\circ})$, and the second centered on the BICEP/Keck field at $(l,b)=(318^{\circ},-61^{\circ})$.
Figure 2: Dust intensity at 353GHz at $(l,b)=(180^{\circ},-10^{\circ})$ (left) and $(l,b)=(318^{\circ},-61^{\circ})$ (right) with an angular resolution of $4.94\arcmin$.
Figure 3: Polarized dust intensity at 353GHz centered at $(l,b)=(180^{\circ},-10^{\circ})$ (left) and $(l,b)=(318^{\circ},-61^{\circ})$ (right), smoothed to $30\arcmin$.
We integrate the dust models in the Planck passband (Planck Collaboration IX, 2014). For the comparison between total intensity maps, we subtract a Wiener-filtered CMB temperature anisotropy map from SMICA from the Planck map.
Intensity and polarization at 353 GHz are displayed in Figures 2 and 3. In total intensity, model d9 filters more of the real data and generates more random small-scale structures than models d1 and d12. We can visually appreciate the impact of this choice in the left panel of Figure 2, near the Galactic plane.
Note that the color scale is different for several of the models in the right panel of Fig. 2. This is because the models have different zero points, mainly due to uncertainty in the CIB monopole. For the generation of model d12, a monopole of 0.09 MJy/sr is subtracted from the DR2 GNILC intensity map. In the case of model d1, it is due to the difference in the CIB monopole between PR2 and PR3 products, as model d1 is based on PR2 data.
@@ 5.1 Maps
5.2 Power spectra
In this section we discuss the power spectrum validation of the PySM dust and synchrotron models. After introducing the methodology in Section 5.2.1, we examine dust and synchrotron emission over large sky areas in Sections 5.2.2 and 5.2.3, respectively, and finally analyze the BICEP/Keck patch specifically in Section 5.2.4.
5.2.1 Methodology
@@
After masking, we compute the power spectra using the anafast function from healpy [^13] https://healpy.readthedocs.io (Zonca et al., 2019) and account for the masking effects by dividing the spectra by $f_{\rm sky}$, the second moment of the mask. We do not implement further corrections for mode-mixing caused by masking, as methods for such corrections assume that the field is free from any inherent mode-coupling (e.g., Hivon et al., 2002), which holds for the CMB but not for the highly non-stationary Galactic emission.
We bin the power spectra into dynamic $\ell$-intervals optimized for map noise and sky fraction. Throughout this work, we present the results as $\mathcal{D}_{\ell}\equiv\ell(\ell+1)C_{\ell}/2\pi$, where $C_{\ell}$ is the power spectrum, with all values expressed in $\mu{\rm K^{2}_{CMB}}$ units.
@@
5.2.2 Dust Emission Over the Sky
To evaluate the performance of the PySM dust models d1, d9, d10, and d12 against real data across a wide range of scales, we present their power spectra computed using both large-area and small-area masks at 353 GHz, where the dust emission is dominant. First, we generate PySM dust model maps at a monochromatic frequency of 353 GHz and smooth them with a $4.82\arcmin$ Gaussian beam to match the resolution of the Planck 353 GHz channel. We color-correct the Planck NPIPE 353 GHz channel maps [^14] https://portal.nersc.gov/project/cmb/planck2020/ (Planck Collaboration et al., 2020) to the same single frequency using a scaling factor 1/1.098 (Planck Collaboration XI, 2020), and subtract the CMB dipole from the NPIPE temperature maps. We then apply identical masks (defined below) to the mean-subtracted PySM dust models and NPIPE A/B detector-split maps. Finally, we compute the auto-spectra for the model maps and the cross-spectra for the NPIPE data splits, as cross-spectra minimize noise bias in the observational data.
@@ 5.2.2 Dust Emission Over the Sky
In Figure 4, we compare the $TT$, $EE$, and $BB$ power spectra of the d1, d9 and d12 models against the NPIPE map. Since the d10 model is identical to the d9 model at 353 GHz, it is not shown separately. Error bars on the cross-spectra are derived from 200 NPIPE detector-split simulations.
Figure 4: The 353 GHz $TT$, $EE$ and $BB$ power spectra for the d1 (dashed lines), d9 (solid lines) and d12 (dotted lines) models, along with the NPIPE detector-split maps (circles), computed using the GAL020, GAL040, GAL060 and GAL080 Galactic masks. Each comparison set is colored to represent the respective sky fraction $f_{\rm sky}$. The d10 model is not shown as it is identical to d9 at 353 GHz.
@@
In polarization, d12 significantly underestimates both $EE$ and $BB$ spectra across all scales. On the contrary, both d1 and d9 align well with $EE$ data across all scales and sky fractions. d9 demonstrates generally smooth power transitions for both $EE$ and $BB$ spectra, although some ripple artifacts emerge around $\ell\sim 100$ in smaller $f_{\rm sky}$ cases, again caused by the parameter optimization. As $f_{\rm sky}$ decreases the NPIPE $BB$ bandpowers rise at $\ell>200$, exceeding the expected noise contribution as computed from simulations. The origin of this rise is unclear, but if it is not simply a statistical fluctuation, it may be due to unmodeled residual noise or instrumental systematic effects, which become more prominent in smaller sky patches. This behavior is not observed in d9, as its small-scale power is derived from template extrapolation.
@@
Figure 5: The binned $BB$ power spectra from the 353 GHz NPIPE detector-split maps and the dust model maps d1, d9 and d12 in two representative fields. The dashed lines indicate the best-fit of the fixed-index power law ($\mathcal{D}_{\ell}^{BB}=A\,\big{(}l/80\big{)}^{\alpha}$ where $\alpha=-0.54$) to the NPIPE data points, with the fit largely driven by the first two bandpowers.
@@
Figure 6: Scatter plots of the mean of the first two 353 GHz $BB$ bandpowers, $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$, and the mean of the last four $BB$ bandpowers, $\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$, as illustrated in Figure 5. The top panel compares $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$ of d1, d9, and d12 against that of NPIPE, while the bottom panel shows the same comparison for $\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$. Each data point represents the results from a circular sky patch with $|b|>30^{\circ}$. Dashed lines indicate the 1:1 ratio.
@@
For our small-area comparison, we instead follow the method described in Planck Collaboration Int. XXX (2016) to define the masks for power spectrum computation. The full sky is divided into 768 patches with a HEALPix grid with $N_{\text{side}}=8$. At the center of each patch, we construct a circular mask covering 400 deg2, with the edges tapered by a $2^{\circ}$ FWHM Gaussian, yielding $f_{\rm sky}\sim 0.008$. Figure 5 presents representative results from two selected circular fields located at Galactic latitudes $|b|>30^{\circ}$ in the northern and southern Galactic hemispheres. Since the computational cost of running simulations for cross-spectra error calculations in each small patch is high, the errors are estimated using analytical approximations of the cross-correlation matrix of power spectra (Tristram et al., 2005).
@@
To assess whether the PySM models can accurately replicate the observed dust $BB$ power spectrum amplitude and scaling in $\ell$, particularly in small, high-Galactic latitude fields relevant to future CMB observations, we calculate the $-\ell$ averaged bandpower $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$ (over $11040\leq\ell<$ 110) and the $-\ell$ averaged bandpower $\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$ (over $370110\leq\ell<$ 370) for both the model and NPIPE maps in each small field with $|b|>30^{\circ}$. These metrics reduce the impact of noise fluctuation in the NPIPE $BB$ data in the circular sky patches, serving as proxies for comparing dust amplitudes between the model and the data at two different angular scales. The results are presented as scatter plots in Figure 6.
@@
Both the d1 and d9 models exhibit power spectral amplitudes that are generally consistent with NPIPE, but d9 demonstrates a significant improvement in correlation, particularly increasing the correlation coefficient from 0.887 to 0.998 in the $-\ell$ regime. This improvement stems from the combination of GNILC large-scale templates and small-scale modulation, which is especially evident in regions with a high signal-to-noise ratio, such as in the $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$ comparison or in small fields with substantial dust amplitudes.
@@
However, in fields where $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$ and $\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$ are smaller than 10 $\mu\rm{K^{2}_{CMB}}$, d9 systematically overestimates the dust amplitude. To investigate the cause, we repeat the same analysis using GNILC template maps and find that the results closely follow the distribution of the d9 scatter points at $-\ell$ in Figure 6. This suggests that the overestimation is likely due to a bias present in regions of the GNILC dust maps with low foregrounds, which propagates from large to small scales through the extrapolated power spectrum fit during our model construction. Conversely, d12 underestimates the dust amplitude in 94% (85%) of the small fields for $-\ell$ ($-\ell$). These opposing trends in d9 and d12 are also observed in individual small fields measured by ongoing $B$-mode experiments, such as the BICEP/Keck field, which will be discussed in Section 5.2.4, and in the SPIDER field independently analyzed by Ade et al. (2025), although the latter results are presented at 150 GHz.
@@
For the comparison of scaling in $\ell$, we introduce the ratio $\mathcal{R}\equiv\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}\Big{/}\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$ as another metric to describe changes in spatial power across the modulation scale. According to this definition, the fixed-index power law $\mathcal{D}_{\ell}^{BB}\propto\ell^{-0.54}$, derived from the analysis of a larger sky region with $f_{\rm sky}=0.8$ by Planck Collaboration XI (2020), yields a value of $\mathcal{R}=1.83$. This estimate is closely aligned with the small-field NPIPE data, which give a value of $\mathcal{R}=1.85\pm 0.93$. The d1, d9, and d12 models instead produce slightly higher, though still consistent, values: $\mathcal{R}=2.03\pm 0.72$, $\mathcal{R}=2.35\pm 0.77$ and $\mathcal{R}=2.26\pm 0.91$ respectively. While the injected small scales in d9 are also generated using an index of $\alpha=-0.54$ (Table 1), this fit is performed in the $bb$ spectrum.
@@
During the development of the PySM models presented in the present study, we used the ratio $\mathcal{R}$ to evaluate the smoothness of the transition in bandpower from large to small scales in small regions. This approach ultimately guided us to adopt the improved modulation map construction method discussed in Section 3.1.2, ensuring a uniform transition between scales. The remaining discrepancy in d9 can be attributed to the non-linear transformation between $D_{\ell}^{bb}$ and $D_{\ell}^{BB}$.
5.2.3 Synchrotron Emission Over the Sky

### v2

@@ IV.2 CO Models
co1
Single line emissions at
115, 230, 346 GHz
Planck PR2 Type-1 maps
smoothed to $1^{\circ}$
co2
"+ $0.1\%$ polarized
co3 " "
Simulated high
Galactic clouds
@@
Galactic CO line emission is strong enough that the $J=1\rightarrow 0$, $J=2\rightarrow 1$, $J=3\rightarrow 2$ transitions have been detected even in the broad photometric Planck bands (planck2013-p03a; planck2014-a12). CO emission can be polarized (Goldreich:1981) and so may be a relevant foreground for CMB polarization analyses (Puglisi:2017). In this section, we describe the implementation of three CO models into the PySM 3 framework. These models are summarized in Table 5.
Because of the intrinsically low degree of CO polarization and the long integration time required to achieve a significant detection, it is difficult to carry out CO polarization surveys, and wide area measurements have not yet been made. In the absence of data-based templates, the approach to modeling CO has been to assume a small degree of polarization applied to total intensity maps. For instance, Puglisi:2017 presented a model to simulate the polarized emission of CO lines in molecular clouds at high Galactic latitudes by considering the 3D spatial distribution of CO in the Galaxy.
All CO models implemented here are based on the CO $J=1\rightarrow 0$, $J=2\rightarrow 1$, $J=3\rightarrow 2$ Type-1 intensity maps [^12] HFI_CompMap_CO-Type1_2048_R2.00.fits derived by planck2013-p03a from Planck data. The CO maps are obtained exploiting the mismatches in the detector bandpasses to recover the CO from the CMB and the other Galactic foregrounds using the Modified Internal Linear Combination Algorithm (MILCA). The templates produced from this analysis are preferred to direct measurements from existing spectroscopic surveys (e.g., Dame:2001) because they are not limited to low Galactic latitudes and have a uniform angular resolution over the sky.
planck2013-p03a also produced Type-2 maps by using single-channel maps at multiple frequencies to separate the CO emission from the astrophysical and CMB signals. However, this method is prone to foreground mis-modeling errors and to instrumental systematics, and the resulting maps contain known discrepancies with other CO observations (planck2014-a12). We thus prefer the Type-1 templates for our purposes.
@@
The second CO model co2 introduces a simple implementation of polarization. The $I$ maps of the three transitions used in co1 are converted to $P$ maps using a sky-constant, user-defined polarization fraction $p$, where the default is p=0.1p=0.1%. The $Q$ and $U$ maps are made from the $P$ maps using the polarization angle of the thermal dust emission as determined by a component separation analysis with Commander (planck2014-a12).
Finally, the third CO model co3 not only accounts for the polarized emission as in co2 but also includes small-scale emission from high-Galactic latitude clouds. We perform a dedicated simulation with the LogSpiral model from Puglisi:2017, which provided the best fit to Planck data, to produce maps of the contribution to the total and polarized CO intensity at sub-degree scales from molecular clouds at high Galactic latitudes. The $I$, $Q$, and $U$ maps generated by the simulation are then added to the $I$, $Q$, and $U$ maps of co2 to produce the final co3 model. This injection introduces CO emission at scales $\lesssim 1^{\circ}$ in regions where the CO templates have very low signal-to-noise ratio.
V Validation and Characterization of Models
In this section, we compare the new foreground models presented in this work both to data and to previous models. In Section V.1, map-based comparisons are made to visually highlight differences between the models and observational data. In Section V.2, we compare power spectra of the dust and synchrotron models with observations. After developing the methodology in Section V.2.1, we demonstrate that the two-point statistics of the stochastic small scales are properly modulated for different regions of sky defined by Galactic masks of varying size. Section V.2.2 compares the dust models to Planck NPIPE maps at 353 GHz. Section V.2.3 compares the PySM synchrotron models with the synchrotron map from the BeyondPlanck analysis for intensity, and with the Planck Revisited reprocessed maps for polarization. Additional validation is performed in the sky patch observed by the BICEP/Keck telescopes in Section V.2.4. Section V.3 compares the level of frequency decorrelation in dust $BB$ emission across the models, demonstrating that all models respect current constraints from Planck observations. Section V.4 quantifies the level of extragalactic contamination in the maps. In Section V.5, we assess the level of non-Gaussianity introduced by the polarization fraction tensor formalism.
@@ V Validation and Characterization of Models
V.1 Maps
We provide several map-level comparisons between PySM dust models and observations. Specifically, we compare data from the Planck third data release PR3 planck2016-l03 and PySM dust models d1, which was widely used as a reference in previous versions of the PySM, and is described in Thorne:2017, d9, which is similar to d1 but uses different input templates and includes small-scale fluctuations of template maps and spectral parameters, as described in Sec. III, and d12, which is conceptually quite different from the other two models. The goal is to provide a qualitative comparison of the spatial characteristics of the models at small scales. We select $16.7^{\circ}\times 16.7^{\circ}$ patches of the sky, and compare the data and models at 353 GHz in both total and polarized intensity. We focus on two patches, one close to the Galactic plane with $(l,b)=(180^{\circ},-10^{\circ})$, and the second centered on the BICEP/Keck field at $(l,b)=(318^{\circ},-61^{\circ})$.
Figure 2: Patches $(16.7^{\circ}\times 16.7^{\circ})$ of dust intensity at 353 GHz centered at $(l,b)=(180^{\circ},-10^{\circ})$ (left, “low latitude”) and $(l,b)=(318^{\circ},-61^{\circ})$ (right, “high latitude” that is centered on the BICEP/Keck field) with an angular resolution of $4.94\arcmin$ for the d1, d9, and d12 dust models and Planck PR3 data.
Figure 3: The same sky regions as Figure 2 shown in 353 GHz polarized dust intensity for the d1, d9, and d12 dust models and Planck PR3 data.
We integrate the dust models in the Planck passband (planck2013-p03d). For the comparison between total intensity maps, we subtract a Wiener-filtered CMB temperature anisotropy map from SMICA from the Planck map.
Intensity and polarization at 353 GHz are displayed in Figures 2 and 3. In total intensity, model d9 filters more of the real data and generates more random small-scale structures than models d1 and d12. The d12 model preserves the observed emission deconvolved from the instrumental beam up to 5′  in intensity (see Section IV.1) whereas d9 preserves the dust template only at scales $\ell<\ell_{1}=100$, corresponding approximately to $1.8^{\circ}$. This is especially evident in high signal-to-noise regions near the Galactic plane, as shown in the left panel of Figure 2.
Note that the color scale is different for several of the models in the right panel of Figure 2. This is because the models have different zero points, mainly due to uncertainty in the CIB monopole. For the generation of model d12, a monopole of 0.09 MJy/sr is subtracted from the DR2 GNILC intensity map. In the case of model d1, it is due to the difference in the CIB monopole between PR2 and PR3 products, as model d1 is based on PR2 data.
@@ V.1 Maps
V.2 Power spectra
In this section we discuss the power spectrum validation of the PySM dust and synchrotron models. After introducing the methodology in Section V.2.1, we examine dust and synchrotron emission over large sky areas in Sections V.2.2 and V.2.3, respectively, and finally analyze the BICEP/Keck patch specifically in Section V.2.4.
V.2.1 Methodology
@@
After masking, we compute the power spectra using the anafast function from healpy [^13] https://healpy.readthedocs.io (Zonca:2019) and account for the masking effects by dividing the spectra by $f_{\rm sky}$, the second moment of the mask. We do not implement further corrections for mode-mixing caused by masking, as methods for such corrections assume that the field is free from any inherent mode-coupling (e.g., Hivon:2002), which holds for the CMB but not for the highly non-stationary Galactic emission.
We bin the power spectra into dynamic $\ell$-intervals optimized for map noise and sky fraction. Throughout this work, we present the results as $\mathcal{D}_{\ell}\equiv\ell(\ell+1)C_{\ell}/2\pi$, where $C_{\ell}$ is the power spectrum, with all values expressed in $\mu{\rm K^{2}_{CMB}}$ units. All power spectrum comparisons presented here are at the native resolution of the data they are compared with.
@@
V.2.2 Dust Emission Over the Sky
To evaluate the performance of the PySM dust models d1, d9, d10, and d12 against real data across a wide range of scales, we present their power spectra computed using both large-area and small-area masks at 353 GHz, where the dust emission is dominant. First, we generate PySM dust model maps at a monochromatic frequency of 353 GHz and smooth them with a $4.82\arcmin$ Gaussian beam to match the resolution of the Planck 353 GHz channel. We color-correct the Planck NPIPE 353 GHz channel maps [^14] https://portal.nersc.gov/project/cmb/planck2020/ (PlanckCollaboration:2020) to the same single frequency using a scaling factor 1/1.098 (planck2016-l11A), and subtract the CMB dipole from the NPIPE temperature maps. We then apply identical masks (defined below) to the mean-subtracted PySM dust models and NPIPE A/B detector-split maps. Finally, we compute the auto-spectra for the model maps and the cross-spectra for the NPIPE data splits, as cross-spectra minimize noise bias in the observational data.
@@ V.2.2 Dust Emission Over the Sky
In Figure 4, we compare the $TT$, EEEE, and $BB$ power spectra of the d1, d9 and d12 models against the NPIPE map. Since the d10 model is identical to the d9 model at 353 GHz, it is not shown separately. Error bars on the cross-spectra are derived from 200 NPIPE detector-split simulations.
Figure 4: The 353 GHz TTTT, EEEE and $BB$ power spectra for the d1 (dashed lines), d9 (solid lines) and d12 (dotted lines) dust models, along with the NPIPE detector-split maps (circles), computed using the GAL020, GAL040, GAL060 and GAL080 Galactic masks. Each comparison set is colored to represent the respective sky fraction $f_{\rm sky}$. The d10 model is not shown as it is identical to d9 at 353 GHz.
@@
In polarization, d12 significantly underestimates both $EE$ and $BB$ spectra, in particular on small scales. This may be at least partly due to the fact that independent random fluctuations are introduced in each layer on small scales, while the large-scale emission is correlated between layers. The independent fluctuations suffer line-of-sight depolarization in the sum of the layers, so that the small scale power is reduced relative to the large-scale power. On the contrary, both d1 and d9 align well with $EE$ data across all scales and sky fractions. d9 demonstrates generally smooth power transitions for both $EE$ and $BB$ spectra, although some ripple artifacts emerge around $\ell\sim 100$ in smaller $f_{\rm sky}$ cases, again caused by the parameter optimization. As $f_{\rm sky}$ decreases the NPIPE $BB$ bandpowers rise at $\ell>200$, exceeding the expected noise contribution as computed from simulations. The origin of this rise is unclear, but if it is not simply a statistical fluctuation, it may be due to unmodeled residual noise or instrumental systematic effects, which become more prominent in smaller sky patches. This behavior is not observed in d9, as its small-scale power is derived from template extrapolation.
@@
Figure 5: The binned BBBB power spectra from the 353 GHz NPIPE detector-split maps and the dust model maps d1, d9 and d12 in two representative fields. The dashed lines indicate the best-fit of the fixed-index power law ($\mathcal{D}_{\ell}^{BB}=A\,\big(l/80\big)^{\alpha}$ where $\alpha=-0.54$) to the NPIPE data points, with the fit largely driven by the first two bandpowers.
@@
Figure 6: Scatter plots of the mean of the first two 353 GHz BBBB bandpowers, $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$, and the mean of the last four BBBB bandpowers, $\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$, as illustrated in Figure 5. The top panel compares $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$ of d1, d9, and d12 against that of NPIPE, while the bottom panel shows the same comparison for $\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$. Each data point represents the results from a circular sky patch with $|b|>30^{\circ}$. Dashed lines indicate the 1:1 ratio.
@@
For our small-area comparison, we instead follow the method described in planck2014-XXX to define the masks for power spectrum computation. The full sky is divided into 768 patches with a HEALPix grid with $N_{\text{side}}=8$. At the center of each patch, we construct a circular mask covering 400 deg2, with the edges tapered by a $2^{\circ}$ FWHM Gaussian, yielding $f_{\rm sky}\sim 0.008$. Figure 5 presents representative results from two selected circular fields located at Galactic latitudes $|b|>30^{\circ}$ in the northern and southern Galactic hemispheres. Since the computational cost of running simulations for cross-spectra error calculations in each small patch is high, the errors are estimated using analytical approximations of the cross-correlation matrix of power spectra (Tristram:2005).
@@
To assess whether the PySM models can accurately replicate the observed dust $BB$ power spectrum amplitude and scaling in $\ell$, particularly in small, high-Galactic latitude fields relevant to future CMB observations, we calculate the low-$\ell$ averaged bandpower $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$ (over $40\leq\ell<110$) and the high-$\ell$ averaged bandpower $\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$ (over $110\leq\ell<370$) for both the model and NPIPE maps in each small field with $|b|>30^{\circ}$. These metrics reduce the impact of noise fluctuation in the NPIPE $BB$ data in the circular sky patches, serving as proxies for comparing dust amplitudes between the model and the data at two different angular scales. The results are presented as scatter plots in Figure 6.
@@
Both the d1 and d9 models exhibit power spectral amplitudes that are generally consistent with NPIPE, but d9 demonstrates a significant improvement in correlation, particularly increasing the correlation coefficient from 0.887 to 0.998 in the low-$\ell$ regime. This improvement stems from the combination of GNILC large-scale templates and small-scale modulation, which is especially evident in regions with a high signal-to-noise ratio, such as in the $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$ comparison or in small fields with substantial dust amplitudes.
@@
However, in fields where $\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}$ and $\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$ are smaller than 10 $\mu\rm{K^{2}_{CMB}}$, d9 systematically overestimates the dust amplitude. To investigate the cause, we repeat the same analysis using GNILC template maps and find that the results closely follow the distribution of the d9 scatter points at low-$\ell$ in Figure 6. This suggests that the overestimation is likely due to a bias present in regions of the GNILC dust maps with low foregrounds, which propagates from large to small scales through the extrapolated power spectrum fit during our model construction. Conversely, d12 underestimates the dust amplitude in 94% (85%) of the small fields for low-$\ell$ (high-$\ell$). These opposing trends in d9 and d12 are also observed in individual small fields measured by ongoing $B$-mode experiments, such as the BICEP/Keck field, which will be discussed in Section V.2.4, and in the SPIDER field independently analyzed by Ade:2025b, although the latter results are presented at 150 GHz.
@@
For the comparison of scaling in $\ell$, we introduce the ratio $\mathcal{R}\equiv\overline{\mathcal{D}_{\text{low-}\ell}^{BB}}\Big/\overline{\mathcal{D}_{\text{high-}\ell}^{BB}}$ as another metric to describe changes in spatial power across the modulation scale. According to this definition, the fixed-index power law $\mathcal{D}_{\ell}^{BB}\propto\ell^{-0.54}$, derived from the analysis of a larger sky region with $f_{\rm sky}=0.8$ by planck2016-l11A, yields a value of $\mathcal{R}=1.83$. This estimate is closely aligned with the small-field NPIPE data, which give a value of $\mathcal{R}=1.85\pm 0.93$. The d1, d9, and d12 models instead produce slightly higher, though still consistent, values: $\mathcal{R}=2.03\pm 0.72$, $\mathcal{R}=2.35\pm 0.77$ and $\mathcal{R}=2.26\pm 0.91$ respectively. While the injected small scales in d9 are also generated using an index of $\alpha=-0.54$ (Table 1), this fit is performed in the $bb$ spectrum.
@@
During the development of the PySM models presented in the present study, we used the ratio $\mathcal{R}$ to evaluate the smoothness of the transition in bandpower from large to small scales in small regions. This approach ultimately guided us to adopt the improved modulation map construction method discussed in Section III.1.2, ensuring a uniform transition between scales. The remaining discrepancy in d9 can be attributed to the non-linear transformation between $D_{\ell}^{bb}$ and $D_{\ell}^{BB}$.
V.2.3 Synchrotron Emission Over the Sky